10. Banyan Co.’s common stock currently sells for $53.25 per share. The growth rate is a constant 5%, and the company has an expected dividend yield of 6%. The expected long-run dividend payout ratio is 50%, and the expected return on equity (ROE) is 10.0%. New stock can be sold to the public at the current price, but a flotation cost of 10% would be incurred. What would be the cost of new equity? Do not round intermediate calculations. Round your answer to two decimal places.

  %
Answer #1

Dividend = yield*price = 0.06*53.25 = 3.195

Growth rate=ROE*retention ratio
growth rate=10*0.5
growth rate = 5
As per DDM
Price(1-flotation %)= Dividend in 1 year/(cost of equity - growth rate)
53.25*(1-0.1) = 3.195/ (Cost of equity - 0.05)
Cost of equity% = 11.67